# CryoNet.Refine: A One-step Diffusion Model for Rapid Refinement of Structural Models with Cryo-EM Density Map Restraints

### v1

@@ 6 Conclusion
In this work, we present CryoNet.Refine, a pioneering one-step diffusion module for cryo-EM atomic model refinement. It critically featuring a novel differentiable density loss and a comprehensive geometry metric loss, not only achieves unparalleled refinement but also opens new avenues: the density loss can be broadly applied to AI-based atomic model building , while the geometry loss offers a powerful guidance of standard geometry for protein structure prediction. CryoNet.Refine thus marks a significant advance for the cryo-EM field. Looking ahead, we aim to extend the specialized geometry loss for DNA/RNA structures, while further enriching the geometry loss family by incorporating a steric clash loss, and demonstrate superior performance on challenging low-resolution density maps, ultimately contributing more profoundly to the entire cryo-EM community.
@@ Appendix A Notation
$\vec{x}=[x^{0},x^{1},x^{2}]$ All-atom coordinates of refined atomic models
$d=[\vec{m},\rho]$ Synthetic density map
$\vec{m}=[\lambda,\mu,\nu]$ Coordinates to locate inside density maps
$\rho$ density values
@@ Appendix C One-step diffusion module
Classical diffusion models define a forward noising process and a multi-step reverse denoising chain (Ho et al., 2020; Song et al., 2021). While powerful, such iterative schemes incur substantial computational cost. Inspired by accelerated sampling approaches such as consistency models (Song et al., 2023) and diffusion distillation (Meng et al., 2022), we reformulated diffusion into a deterministic one-step refinement tailored for cryo-EM structural optimization.
@@
Thus, each recycle performs exactly one deterministic refinement step, in contrast to the multi-step stochastic denoising trajectories used in classical diffusion models.
@@
4:Refined structure $\mathbf{x}_{n}$
5:for recycle $i=1,2,\ldots,n$ do
6:Initialize $\mathbf{x}\leftarrow\mathbf{x}_{0}$ $\triangleright$ Current structure estimate
@@
11:Step 3: Perform structure update (One-step)
@@
13:Step 4: Compute loss and backpropagate
@@
17:return $\mathbf{x}$ $\triangleright$ Final refined structure after $n$ recycles
@@ Appendix D Density Generator
Our current density generator is not a neural network but a physics-based simulator. It constructs the target density volume by summing Gaussian functions centered on atomic positions with the simulation process specifically tailored to the target resolution. This procedure conceptually follows the molmap implementation in UCSF ChimeraX(Pettersen et al., 2021).
@@
To evaluate how well the simulated maps agree with input experimental maps, we calculate their correlation coefficients across all cases. The average correlation is 0.749, with over 80% of all cases exceeding 0.70. Nevertheless, simulated maps inevitably fall short of capturing the complex characteristics introduced by diverse experimental instruments and imaging conditions. In particular, secondary structure elements, map artifacts, and noise patterns in our simulated densities differ from those observed in experimental maps. To address these limitations, future work will explore replacing the simulation-based density generator with a deep learning–based generative model, which could better reproduce the statistical and structural properties of experimental cryo-EM maps and yield a more informative and accurate density loss.
@@ F.1 Recycle Strategy
The Recycling strategy of CryoNet.Refine is setting 300 iterations as maximum with an early-stop mechanism. To demonstrate the rationale behind this, we picked a protein structure from our dataset (PDB-6ksw) and trace its model-map correlation coefficients (CC) values after each recycle iteration, as shown in Figure 12.
These 4 CC values basically follow a similar pattern. The early-stop mechanism effectively spots a good termination where the curves converge after the uphill slope.
F.2 Diffusion Sampling Steps

### v2

@@ 6 Conclusion
In this work, we present CryoNet.Refine, a pioneering one-step diffusion module for cryo-EM atomic model refinement. It critically features a novel differentiable density loss and a comprehensive geometry metric loss, achieving unparalleled refinement accuracy.
Our framework opens new avenues for the field: the density loss can be broadly applied to AI-based atomic model building, while the geometry loss offers essential guidance for protein structure prediction. Specifically, these differentiable constraints are model-agnostic and can complement the implicit priors of foundation models like AlphaFold3. By integrating them into conditional diffusion sampling strategies, future research could bridge the gap between generative priors and the atomic-level stereochemical precision required for high-resolution structures.
CryoNet.Refine thus marks a significant advance for the cryo-EM community. Looking ahead, we aim to extend the specialized geometry loss for DNA/RNA structures and incorporate a steric clash loss. We also acknowledge the computational cost associated with per-structure optimization; therefore, future work will focus on enhancing efficiency through parallel refinement frameworks and faster convergence strategies. Ultimately, we intend to demonstrate superior performance on challenging low-resolution density maps, establishing CryoNet.Refine as an essential tool for structural biology.
@@ Appendix A Notation
$\vec{\mathbf{x}}=[\mathbf{x}^{0},\mathbf{x}^{1},\mathbf{x}^{2}]$ All-atom coordinates of refined atomic models
$d=[\vec{{\bm{m}}},\bm{\rho}]$ Synthetic density map
$\vec{{\bm{m}}}=[\bm{\lambda},\bm{\mu},\bm{\nu}]$ Grid points of density maps
$\hat{\bm{\rho}}$ Synthetic density values
$\bm{\rho}$ Density values retrieved from the input density map
@@ Appendix C One-step diffusion module
Classical diffusion models define a forward noising process and a multi-step reverse denoising chain (Ho et al., 2020; Song et al., 2020). While powerful, such iterative schemes incur substantial computational cost. Inspired by accelerated sampling approaches such as consistency models (Song et al., 2023) and diffusion distillation (Meng et al., 2022), we reformulated diffusion into a deterministic one-step refinement tailored for cryo-EM structural optimization.
@@
4:Refined structure $\mathbf{x}_{\text{final}}$
5:Initialize model parameters $\theta$
6:for recycle $i=1,2,\ldots,n$ do
@@
10:Step 2: Perform structure update (One-step)
@@
12:Step 3: Compute loss and backpropagate
@@
17:return $\mathbf{x}_{\text{pred}}$
@@
It is important to distinguish between the diffusion sampling mechanism and the workflow. The term “One-step” refers to the efficiency of our diffusion sampler (Eq. 13). Unlike stochastic samplers that require hundreds of iterative denoising steps to generate a structure, our preconditioned module maps the initial structure $\mathbf{x}_{0}$ directly to the predicted refined state in a single forward pass.
In contrast, “Recycling” refers to the outer test-time optimization loop. Although the sampler is capable of generating a structure in one step, achieving high-precision fit to experimental density requires iterative tuning. In each recycle, we perform a one-step forward pass using the fixed input $\mathbf{x}_{0}$, compute density and geometry losses on the output, and backpropagate gradients to update the model parameters $\theta$. This iterative optimization allows the diffusion module to progressively customize its weights to the specific instance, minimizing the loss over multiple cycles.
C.1 Architectural Differences from AlphaFold3
While CryoNet.Refine initializes its parameters from the weights of Boltz-2, its architectural design and inference workflow diverge fundamentally from the standard AlphaFold3 framework to address the specific challenges of cryo-EM refinement. The key distinctions are summarized below:
  1. 1.
    Input Initialization: Standard AlphaFold3 is a generative model designed for de novo prediction, initializing the diffusion process from Gaussian noise ($\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$) to generate diverse conformations. In contrast, CryoNet.Refine is a refinement framework. We initialize directly from the initial atomic model, treating it as a noisy state that requires correction against the density map.
  2. 2.
    Sampling Trajectory: AlphaFold3 employs a stochastic multi-step sampler to traverse the reverse diffusion process, which is computationally expensive and introduces randomness. CryoNet.Refine reformulates this into a one-step prediction, mapping the input state directly to the denoised structure in a single pass for efficiency.
  3. 3.
    Optimization Paradigm: AlphaFold3 typically operates in a fixed-weight inference mode. Conversely, CryoNet.Refine adopts a test-time optimization strategy. During the refinement loop (Algorithm 1), we explicitly compute gradients from the density and geometry losses to update the diffusion module parameters ($\theta$). This allows the network to learn the specific features of the experimental density map for each individual case.
  4. 4.
    Module Adaptations and Trainability: To specialize for density-guided refinement, we streamlined the modular architecture.
    • Multi-sequence alignment (MSA) processing, confidence heads (pLDDT/PAE), and internal architecture-level recycling are removed, as the method relies on physical density constraints rather than evolutionary information.
    • The Sequence Embedder, Atom Encoder, and Pairformer are kept frozen to preserve robust structural priors derived from pre-training. Only the Diffusion Module is trainable, focusing the optimization capacity solely on coordinate adjustment driven by the experimental map.
@@ Appendix D Density Generator
Our current density generator builds the target density volume by summing Gaussian functions centered on atomic positions with the simulation process specifically tailored to the target resolution. This procedure conceptually follows the molmap implementation in UCSF ChimeraX(Pettersen et al., 2021).
A density map $d$ can be discretized as a group of 4-dimensional vectors ${\bm{\lambda}},\bm{{\mu}},\bm{{\nu}},\bm{\rho}]$, where $\vec{{\bm{m}}}=[\bm{\lambda},\bm{\mu},\bm{\nu}]$ act as grid points that can specify a certain location and $\bm{\rho}$ is the density values at each grid point, with the origin of coordinates nearest to $\vec{{\bm{m}}}_{\text{min}}$ and farthest to $\vec{{\bm{m}}}_{\text{max}}$. For computation of the overlapped region, we first locate this region $\forall\vec{{\bm{m}}}_{o}\in\mathcal{S}_{o}$ between the input density map $d_{0}$ and the synthetic density map $d_{i}$ through the coordinate system:
After locating $\mathcal{S}_{o}$, we can select all the density values within it from $d_{0}$ and $d_{i}$ and get their subsets $d_{0\bigcap o}$ and $d_{i\bigcap o}$. Thus, the in this overlapped region $\mathcal{S}_{o}$, our synthetic density values and the input density values will be defined as:
$\hat{\bm{\rho}}=\bm{\rho_{0\bigcap o}},\quad\bm{\rho}=\bm{\rho_{i\bigcap o}}$
For the implementation of synthetic density map generation, we follow the molmap algorithm in ChimeraX(Pettersen et al., 2021), utilizing the all-atom coordinates of the refined atomic model at the $i^{th}$ recycle $\vec{{\bm{x}}}_{i}=[\bm{x}_{i}^{0},\bm{x}_{i}^{1},\bm{x}_{i}^{2}]$ and output a box-shaped density map $d_{i}=(\vec{{\bm{m}}}_{i},\bm{\rho}_{i})$, where $\vec{{\bm{x}}}_{i}\in\mathbb{R}^{3\times N_{\text{atom}}}$ and $\vec{{\bm{m}}}_{i}\in\mathbb{R}^{3\times N_{\text{atom}}\times L}$, noting that the parameter $L$ stands for the number of grid points for each atom per axis, and its value is obtained by:
where resres is the input density map resolution and $v$ represents its voxel size. The key idea here is to construct a Gaussian sphere for each atom, with $r_{\mathcal{G}}$ and $s_{\mathcal{G}}$ respectively meaning Gaussian radius and the radius of a Gaussian sphere. We expect that $\vec{{\bm{m}}}_{i}$ turns out to be a box as small as possible with all-atom coordinates $\vec{{\bm{x}}}_{i}$ inside it as well as a margin of Gaussian sphere around it, so:
$\vec{\bm{m}}_{i_{\text{min}}}=[\lfloor\min(\bm{x}_{i}^{0})\rfloor,\lfloor\min(\bm{x}_{i}^{1})\rfloor,\lfloor\min(\bm{x}_{i}^{2})\rfloor],\quad$ $\vec{\bm{m}}_{i_{\text{max}}}=[\lceil\max(\bm{x}_{i}^{0})\rceil,\lceil\max(\bm{x}_{i}^{1})\rceil,\lceil\max(\bm{x}_{i}^{2})\rceil]$
$\vec{\bm{m}}_{i\_\text{size}}=\lceil(\vec{\bm{m}}_{i_{\text{max}}}-\vec{\bm{m}}_{i_{\text{min}}})/v+2\cdot s_{\mathcal{G}}+1\rceil,\quad$ $\vec{\bm{m}}_{i_{\text{center}}}=(\vec{\bm{m}}_{i_{\text{min}}}+\vec{\bm{m}}_{i_{\text{max}}})/(2\cdot v)$
$\vec{\bm{m}}_{i_{\text{min}}}=\lfloor(\vec{\bm{m}}_{i_{\text{center}}}-\vec{\bm{m}}_{i\_\text{size}})/2\rfloor,\quad$ $\vec{\bm{m}}_{i_{\text{max}}}=\lfloor(\vec{\bm{m}}_{i_{\text{center}}}+\vec{\bm{m}}_{i\_\text{size}})/2\rfloor$
$\text{for }n\in(1,2,...,N_{\text{atom}})\quad{\bm{\rho}}_{\lambda}^{n}=w_{n}\cdot\exp(-\frac{(\bm{x}_{i}^{0})^{n}+{\bm{l}}}{2})\text{, }{\bm{l}}=[-s_{\mathcal{G}},...,0,-s_{\mathcal{G}}],$ (14)
@@
To evaluate how well the simulated maps agree with input experimental maps, we calculate their correlation coefficients across all cases. The average correlation is 0.892, significantly improving from 0.803 when using the USCF ChimeraX molmap method. Nevertheless, simulated maps inevitably fall short of capturing the complex characteristics introduced by diverse experimental instruments and imaging conditions. In particular, secondary structure elements, map artifacts, and noise patterns in our simulated densities differ from those observed in experimental maps. To address these limitations, future work will explore replacing the simulation-based density generator with a deep learning–based generative model, which could better reproduce the statistical and structural properties of experimental cryo-EM maps and yield a more informative and accurate density loss.
@@ F.1 Recycle Strategy
The recycling strategy of CryoNet.Refine sets a maximum of 300 iterations combined with an early-stop mechanism. To demonstrate the rationale behind this, we picked a protein structure from our dataset (PDB-6ksw) and trace its model-map correlation coefficients (CC) values after each recycle iteration, as shown in Figure 12.
This trajectory serves as a sensitivity analysis regarding the number of recycling steps, revealing two distinct phases in the refinement process:
  1. 1.
    High-Sensitivity Phase (Recycles 01000-100): The CC metrics increase sharply, indicating that the model is highly responsive to the initial density-aware gradient updates as it resolves major structural conflicts.
  2. 2.
    Robust Convergence Phase (Recycles >100>100): The curves plateau asymptotically, showing that the refinement becomes robust and insensitive to additional iterations once the structure matches the density map constraints.
These 4 CC values basically follow a similar pattern. Consequently, the setting of 300 iterations provides a sufficient safety margin for convergence, while the early-stop mechanism effectively identifies the transition to the plateau phase, avoiding unnecessary computation.
F.2 Diffusion Sampling Steps and Comparison with Classical Diffusion
@@
To validate our design choice of a one-step diffusion module, we conducted a comparative ablation study against a classical multi-step diffusion approach.
Experimental Setup. For the classical diffusion baseline, we employed a standard stochastic sampling trajectory with $T=200$ steps within one recycle, a setting typical for generative protein structure models like Boltz-2 or AlphaFold3. The inference process started from the initial AlphaFold3 prediction. At each step $t$, the network predicted the denoised structure, and Gaussian noise was injected according to a standard variance schedule before proceeding to step t1t-1.
[…]
Table 4: Quantitative comparison between CryoNet.Refine (One-step) and Classical Diffusion (200 steps) across 27 cases. The multi-step approach fails to preserve structural biology constraints when applied to refinement tasks.
Category Metrics CryoNet.Refine (One-step) Classical Diffusion (200 steps)
Model-map Correlation Coefficients $\rm CC_{mask}\uparrow$ 0.65 0.30
$\rm CC_{box}\uparrow$ 0.58 0.35
$\rm CC_{mc}\uparrow$ 0.66 0.31
$\rm CC_{sc}\uparrow$ 0.63 0.31
$\rm CC_{peaks}\uparrow$ 0.47 0.24
$\rm CC_{volume}\uparrow$ 0.65 0.34
[…]
We suspect that the degradation arises from 2 main factors: (1) We employ a deterministic sampler, which intrinsically requires fewer steps; and (2) The diffusion module receives as input an already complete and physically reasonable macromolecular structure, rather than random noise as in the training regimes of AlphaFold3 or Boltz-2 for de novo structure prediction.
These findings confirm that for the specific task of structural refinement against density maps, a targeted one-step diffusion scheme is far superior to classical multi-step generation.
@@ F.3 Ablation Study
F.4 Comparison with Direct Numerical Optimization
To investigate whether the generative capabilities of our diffusion model are strictly necessary, or if the structural improvements are merely due to the proposed differentiable loss functions, we implemented a pure numerical coordinate optimization baseline. This baseline involves no neural networks, no diffusion process, and no learned structural priors.
Experimental Setup. In this baseline, we treated the atomic coordinates of the initial AlphaFold3 prediction as learnable parameters ($\mathbf{x}\in\mathbb{R}^{3\times N_{\text{atoms}}}$). The optimization process was designed to be as close as possible to the CryoNet.Refine inference loop:
  • Inputs: The initial AlphaFold3 structure and the experimental cryo-EM density map.
  • Objective Function: We utilized the identical composite loss function used in CryoNet.Refine: $\mathcal{L}=\mathcal{L}_{\text{den}}+\mathcal{L}_{\text{geo}}$, employing the exact same weights ($\gamma$) for all density and geometric terms.
  • Optimization: The coordinates were updated directly via backpropagation using a Stochastic Gradient Descent (SGD) optimizer with a momentum of 0.9.
  • Schedule: To ensure a fair comparison with our diffusion model, we set the maximum number of iterations (recycles) to 300, matching the diffusion inference budget. An early stopping mechanism with a patience of 20 iterations was implemented, along with a ReduceLROnPlateau scheduler to dynamically adjust the learning rate.
[…]
This indicates that without the generative and exploratory capabilities of the diffusion model, the gradient descent process gets trapped in local minima close to the initial structure. It satisfies the geometric constraints (which are local) but cannot traverse the energy landscape to find the global conformation that matches the experimental density. Thus, the one-step diffusion module is essential for balancing data fidelity with biological plausibility.
Table 6: Comparison between CryoNet.Refine and Direct Numerical Optimization (SGD) on a subset of 27 protein complexes. The numerical baseline optimizes geometry well but fails to fit the density map.
Category Metrics AlphaFold3 Phenix. real_space_refine Numerical Optimization CryoNet. Refine
Model-map Correlation Coefficients $\rm CC_{mask}\uparrow$ 0.44 0.61 0.46 0.65
$\rm CC_{box}\uparrow$ 0.44 0.55 0.47 0.58
$\rm CC_{mc}\uparrow$ 0.45 0.62 0.46 0.66
$\rm CC_{sc}\uparrow$ 0.44 0.61 0.45 0.64
$\rm CC_{peaks}\uparrow$ 0.31 0.44 0.32 0.47
$\rm CC_{volume}\uparrow$ 0.47 0.61 0.48 0.65
[…]
F.5 Runtime Profiling
To provide deeper insights into the computational efficiency of CryoNet.Refine, we conducted a detailed runtime profile analysis.
We profiled the execution time of a single refinement recycle for a representative protein complex (e.g., PDB-6ksw, $\sim$1,900 residues). The breakdown of time consumption per iteration is detailed in Table 7.
The results indicate that the Gradient Computation (Backpropagation) and the Loss Calculation are the two primary computational bottlenecks. The forward pass of the diffusion model is relatively efficient due to the one-step sampling.
Table 7: Runtime profile of a single refinement recycle (averaged over total recycle).
Component Time (ms) Percentage (%)
Data Loading & Preprocessing 150 $\sim$6%
Model Forward Pass (One-step) 200 $\sim$7%
Loss Calculation (Density + Geometry) 940 $\sim$35%
Backpropagation & Optimization 1420 $\sim$52%
Total per Recycle 2710 100%